OpenStudy (anonymous):

what is 7/8 minus 6/7

OpenStudy (anonymous):

first get common denominators which just multiply 8 by 7 and you get 56 then multiply the top of first by 8 and the 2nd by 8 and you get \[\frac{ 49 }{ 56 } + \frac{ 48 }{ 56 } = \[\frac{ 97 }{ 56 }\] and reduce if needed

OpenStudy (anonymous):

can you simplify it

OpenStudy (anonymous):

no need. already reduced

OpenStudy (anonymous):

unless you want it in a mixed number form

OpenStudy (anonymous):

it has to be on of these 0 one-half 1 one and one-half

OpenStudy (anonymous):

1. 0 2. 1/2 3. 1 4. 1 1/2

OpenStudy (anonymous):

Are you sure you got the problem correct? When done on a calculator \[\frac{ 7 }{ 8 }-\frac{ 6 }{ 7 }=\frac{ 1 }{ 56 }\]